Child pages
  • Engineering Report Day 11 TasMars2013
Skip to end of metadata
Go to start of metadata

Engineering Report: 02/05/2013
Emma Braegen

 
Was Power Interrupted:
no
Time Of Power Outage:

Length Of Power Outage:

Diesel Fuel Level:
0.17
Propane Fuel Level:
0.68
Gasoline Fuel Level:
0.5
Fuel Notes And Comments:

Atv Oil Quarts:
N.A.
Car Oil Quarts:
N.A.
Oil Notes And Comments:

Trailer Water Fill Needed:
no
Trailer Water Refilled:
no
Potable Water Meter Reading:
62666.3
Water Notes And Comments:

Greenhouse Notes And Comments:

Telescope Used:
no
Observatory Notes And Comments:

H A Lp Bugs:

Network Failed:
no
Network Throttled:
yes
Network Signal Strength:
183
Network Download Allowance:
0
Network Notes And Comments:
Internet was throttled again but can't determine what is causing the high internet usage
Hvac Air Conditioner Used:
no
Hvac Furnace Used:
yes
Hvac Lab Wall Heater Used:
no
Hvac Notes And Comments:

Hab Car Used:
no
Hab Car Oil Checked:
no
Hab Car Tire Status:

Hab Car Odometer:

Hab Car Notes And Comments:

Atv One Used:
no
Atv One Oil Checked:
no
Atv One Fuel Used Gals:

Atv One Tires Status:

Atv One Hours Used Day:

Atv One Hours Used Year:

Atv One Notes And Comments:

Atv Two Used:
no
Atv Two Oil Checked:
no
Atv Two Fuel Used Gals:

Atv Two Tires Status:

Atv Two Hours Used Day:

Atv Two Hours Used Year:

Atv Two Notes And Comments:

Atv Three Used:
no
Atv Three Oil Checked:
no
Atv Three Fuel Used Gals:

Atv Three Tires Status:

Atv Three Hours Used Day:

Atv Three Hours Used Year:

Atv Three Notes And Comments:

Atv Four Used:
no
Atv Four Oil Checked:
no
Atv Four Fuel Used Gals:

Atv Four Tires Status:

Atv Four Hours Used Day:

Atv Four Hours Used Year:

Atv Four Notes And Comments:

Atv Five Used:
no
Atv Five Oil Checked:
no
Atv Five Fuel Used Gals:

Atv Five Tires Status:

Atv Five Hours Used Day:

Atv Five Hours Used Year:

Atv Five Notes And Comments:

Comments:

Today I did some work on the dust suit experiments. I collected dust from different areas around the hab and put it into sample containers. I should be able to get these analysed after leaving MDRS. I also tried a first run of experiments with the rest of the crew wearing painters coveralls with the space suits on EVA to see what sort of dust was collected. I had some difficulty working out the best method of collection once people returned from EVA. The idea was to shake off dust onto drop sheets or into bags and weigh changes in the bags and dust suits. Since it's not so dusty it was actually hard to get a decent sample size. I will think about what methods to use and try later in the week to see what works. Hopefully it will get more dusty here to make it easier to get a result.
No engineering issues except for the internet bandwidth still being used up too fast. It's unclear what is causing the problem since we turned off the webcam and don't seem to be using the internet very much.

  • No labels